Find the equation of the line through each given pair of points ( -1,-3), (2,-1)
Answer by unlockmath(1688)

Hello,
Let's first find the slope for the equation. This (-1,-3), (2,-1) is shown as:
-3--1 over -1-2 to give us:
2/3
Now using the y=mx+b where m is slope we can find b:
-3=2/3(-1)+b Solve for b to get:
b=-2 1/3 So the equation is:
y=2/3(x)-2 1/3
Make Sense?
RJ
www.math-unlock.com